### v1

# Inland-LOAM: Voxel-Based Structural Semantic Mapping for Inland Waterways

Abstract
Accurate and up-to-date geospatial information is crucial for enhancing the safety and autonomy of Inland Waterway Transport (IWT). Inland Electronic Navigational Charts (IENC) often lack detailed real-time information, conventional LiDAR-based Simultaneous Localization and Mapping (SLAM) methods have difficulty handling distinctive inland waterway characteristics, such as dynamic shorelines and challenging reflective water surfaces. These factors lead to significant localization drift, particularly in the vertical direction, and produce dense point cloud maps without the semantic richness required for effective autonomous decision making.
This paper introduces Inland-LOAM, a comprehensive LiDAR-based SLAM framework specifically for inland waterway environments. Firstly, we present an improved feature extraction method adapted to the unique geometric characteristics of waterways, combined with a joint optimization framework that incorporates the water surface as a global planar constraint to significantly mitigate vertical drift. Secondly, we propose an innovative pipeline to transform dense 3D point cloud outputs into structured 2D semantic maps by constructing semantic voxel grids and performing geometric analyses (roughness, planarity, and slope). This enables classification into structurally meaningful categories and supports real-time computation of critical navigational parameters such as vertical bridge clearances. Additionally, an automated map conversion module efficiently extracts shoreline boundaries via ray-casting and exports these features into a lightweight, IENC-compatible format.
Extensive evaluations were conducted using a diverse, real-world dataset featuring urban canals, vegetated banks, and bridge crossings. Comparative results demonstrate that Inland-LOAM achieves superior localization accuracy over state-of-the-art LiDAR SLAM methods. Generated semantic maps and extracted shoreline boundaries align with real-world conditions, as verified through comparisons with reference maps, thus providing reliable and up-to-date geospatial information to enhance navigational situational awareness. Both the dataset and the algorithm are publicly available to support future research. Upon publication, the code and datasets will be made publicly accessible.
keywords:
SLAM, Uncrewed surface vehicle, Inland waters, Light detection and ranging (LiDAR), Semantic mapping
[…]
[inst1]organization=Department of Mechanical Engineering,uni=KU Leuven, postcode=3001, city=Leuven, country=Belgium
[…]
[inst2]organization=Department of Computer Science,uni=KU Leuven, postcode=8200, city=Bruges, country=Belgium
[…]
[inst3]organization=Faculty of Mechanical Engineering,uni=TU Eindhoven, postcode=5612 AZ, city=Eindhoven, country=Netherlands
[…]
[inst4]organization=Flanders Make – Leuven,postcode=3001, city=Leuven, country=Belgium
1 Introduction
IWT constitutes an essential component of Europe’s freight infrastructure, spanning a network exceeding 41,000 km, interlinking major cities and industrial hubs across 13 interconnected Member States [1]. As efforts increase to shift freight from congested road and rail networks, the importance of accurate geospatial information and detailed environmental models for managing and navigating these waterways grows [2]. However, navigating these waterways involves safety risks, with studies identifying human error as the primary cause in up to 70% of maritime incidents [3]. In this context, advancements in automation, especially the deployment of uncrewed surface vehicles (USVs), are driving a transformation in inland waterway transport. USVs play a crucial role not only by mitigating the risks associated with human factors, but also by promising improved cargo transport efficiency and facilitating innovative logistics services, including operations within challenging environments such as congested urban canal networks [4]. However, the effective operation of USVs is challenging due to the complex and dynamic nature of inland waterways. Safe and efficient navigation mainly depends on robust systems for perception, localization, and mapping. These systems must enable the vessel to reliably recognize diverse surroundings, including shorelines, infrastructure, vegetation, and overhead obstacles such as bridges, to operate safely.
Traditionally, navigation on inland waterways relies highly on Global Navigation Satellite Systems (GNSS) for positioning, complemented by standard nautical charts like IENC for contextual spatial data [5]. However, GNSS frequently suffer from severe attenuation or complete outages in typical IWT scenarios, particularly during passages under bridges, within narrow urban canals flanked by tall buildings ’urban canyons’, or along waterways with dense vegetation canopies [6, 7]. This documented unreliability highlights the need for onboard localization systems that work independently [8]. Furthermore, existing geospatial data products, IENC, while providing important routing and regulatory data, often lack the spatial detail or real-time updates required for accurate autonomous maneuvering and perception. Representations of shorelines and their surrounding features, including vegetation, shoreline constructions, and other obstacles, are often generalized or outdated in existing maps [9, 10]. A key limitation of IENC is the representation of safety parameters, such as vertical bridge clearance, using static data. This approach fails to account for dynamic water level variations, a significant operational challenge in IWT affecting actual clearances [11]. As a result, relying on static, low-resolution geospatial information hinders the deployment of advanced autonomous functions.
LiDAR-based SLAM has developed as a strong technique in terrestrial robotics and autonomous driving, making it possible to build precise 3D maps and accurate localization even in environments without GNSS [12, 13]. Applying LiDAR SLAM to the IWT is a viable way to overcome the limitations of previous approaches. However, traditional LiDAR SLAM algorithms face particular challenges in this context. The main challenges come from the structure of the environment: unlike urban scenes with abundant planar facades (buildings) and consistent features (lane markings), inland waterways predominantly feature irregular, natural shorelines often covered with vegetation [11]. This lack of structured geometry and standardized visual cues affects the reliable extraction of edges and planar features necessary by algorithms like LOAM [14] and LeGO-LOAM [15]. In addition to these algorithmic challenges, the development and evaluation of SLAM algorithms for IWT is hindered by the lack of suitable benchmark datasets.
Beyond achieving accurate geometric reconstruction, effective environmental understanding demands semantic understanding, e.g. distinguishing navigable water from vegetated banks, identifying a potential docking quay wall, or determining safe passage under a bridge. Standard SLAM methods typically do not provide this layer of structural semantic information. Finally, a gap exists in translating the rich, high-resolution data from SLAM systems into formats aligned with IENC navigation standards.
Figure 1: Our experimental platform, a catamaran named the Maverick, was equipped with a Sensor Box for situational awareness and data collection.
To address these challenges, we propose Inland-LOAM, a novel voxel-based system designed for structural semantic mapping in inland waterway environments. This method achieves robust LiDAR-based localization and mapping adapted for water surfaces, extracts structural semantic information using geometric analysis within a voxel map, and generates enhanced environmental representations suitable for navigation and integration with standard chart formats. The contributions of this paper are as follows:
  1. 1.
    Inland-LOAM is specifically enhanced for robustness in GNSS-denied inland waterway environments through improved feature extraction performance and the incorporation of planar surface optimization.
  2. 2.
    A novel voxel-based structural semantic mapping approach using roughness, planarity, and slope analysis to generate 2D maps and calculating real-time bridge clearance.
  3. 3.
    An automated map conversion method to translate the high-detail structural semantic SLAM map into an IENC-compatible.
  4. 4.
    The proposed framework was demonstrated in a diverse set of real-world inland waterway scenes, including ports, urban canals, vegetated banks, bridges, and lifting gates. The experimental platform used for data collection is shown in Fig. 1. All data are released as part of the publicly available dataset to support further research.
2 Related Work
This section reviews existing literature relevant to LiDAR-based mapping and semantic understanding in the context of inland waterways, focusing on inland mapping methodologies, LiDAR odometry techniques, semantic scene understanding, and map integration with navigation standards.
2.1 Mapping in Inland Waterway Environments
Inland waterways are mapped using various methods depending on their scale and purpose. Large-scale remote sensing, using airborne or satellite platforms with LiDAR, multispectral imagery, or synthetic aperture radar (SAR), provides maps and monitoring of waterway networks and adjacent land cover [16, 17]. However, these top-down approaches often yield static data lacking the resolution and perspective needed for capturing critical navigational details. LiDAR-based onboard mapping offers a promising solution for detailed mapping in inland scenarios [18, 19, 20, 21]. However, deploying LiDAR SLAM in IWT presents distinct complexities compared to established terrestrial applications [22, 23]:
Water surface Interference: Reflections on the surface can distort LiDAR returns, causing noisy or inaccurate measurements. Also, signal absorption along the z-axis weakens vertical constraints [24]. Due to continuous motion induced by waves and currents, vessels cannot be modeled using non-holonomic motion constraints [25].
Long Trajectories and Sparse Loops: The nature of waterway transport involves long traverses with infrequent returns to previously mapped areas, demanding high odometry accuracy due to limited opportunities for loop closure [26].
Data Scarcity: The robust evaluation of SLAM algorithms is hampered by the lack of comprehensive public data sets that capture the various conditions (narrow canals, locks, vegetation, bridges) and operational realities of European IWT; existing datasets like USVInland [27] often have limitations in scope or relevance.
2.2 LiDAR Odometry and Mapping and Variants
LOAM [14] pioneered efficient, real-time LiDAR odometry based on the roughness of each point on its respective ring. Its low drift and computational efficiency made it has been ranking top-3 in the KITTI odometry leaderboard [28]. LeGO-LOAM [15] built upon LOAM by incorporating point cloud segmentation to separate ground points from other objects and clustering non-ground points. This allows focusing feature extraction on more salient objects and significantly reduces computational cost, making it suitable for platforms with limited processing power. Subsequent works, such as F-LOAM [29], further focused on developing lightweight LOAM variants with optimized efficiency. For flat-ground scenarios, HDL-Graph-SLAM [30] employs graph optimization using NDT scan matching integrated with ground plane constraints. Ground-SLAM [31] specifically addresses structured multi-floor environments by extracting ground segments and modeling them as planar landmarks via Closest Point (CP) parameterization. Recently, KISS-ICP [32] achieved high performance through simple but effective point-to-point registration and accurate motion compensation. However, methods explicitly incorporating ground plane constraints typically assume the existence of a continuous solid ground. Additionally, ICP-based approaches, despite their computational efficiency, often overlook the richer geometric information provided by feature-based methods.
2.3 Semantic SLAM and Scene Understanding
Semantic SLAM commonly relies on Deep Learning (DL), SuMa++ [33] use Fully Convolutional Networks (FCNs) [34] for segmentation that allows filtering of dynamic objects. Others [35] leverage Transformers [36] for integrated feature extraction and segmentation, enabling selective processing focused on static scene components to enhance localization accuracy. While effective for object recognition, such DL methods typically rely on large labeled datasets.
In contrast, geometric-based approaches derive semantic understanding without extensive annotations. $\mathcal{PLC}$-LiSLAM [37] detects geometric primitives, such as planes, lines, and cylinders, as stable landmarks for SLAM optimization. Similarly, [38] achieves forest localization by clustering to identify trees and matching them to aerial canopy maps. These methods leverage inherent geometric structures for environmental understanding, avoiding the need for learned object appearance models.
2.4 Inland Navigation Standards and Map Integration
IENCs are vital for navigation on European inland waterways, delivering essential details on waterway layouts, aids to navigation, locks, and bridges [39]. Recent studies have explored enhancing IENCs with sensor data. For example, Van Baelen et al. [9] used onboard LiDAR to enhance situational awareness for inland vessels, detecting dynamic environments via a KDTree approach and aligning them with IENC data. Pieper and Hahn [10] proposed a 3D LiDAR sensor fusion method to improve ENC accuracy in harbor areas. Our prior work, Luo [18], dynamically updated IENCs using clustering techniques on LiDAR data. While these efforts highlight the potential of LiDAR for IENC updates, they often depend on direct projection and ad-hoc feature fitting. Similarly, Fredriksson et al. [40] converted voxel maps into 2D occupancy grids for path planning by projecting verified free space and preserving height and slope metadata, suggesting a promising pathway for efficient 3D-to-2D map conversion. However, existing research either validates isolated ENC objects or compresses 3D maps for robotic planning, lacking an end-to-end algorithm to transform LiDAR-derived structural and semantic data into standardized geospatial overlays suitable for IENCs.
3 Methodology
3.1 System Overview
The Inland-LOAM framework, as shown in Fig. 2, is designed to build structural semantic maps of inland waterways for enhanced situational awareness, only relying on LiDAR. The system comprises two primary modules: first, a robust LiDAR Odometry and Mapping (LOAM) pipeline for accurate pose estimation and initial environmental representation; and secondly, a semantic interpretation and conversion module to produce enhanced, IENC-compatible maps.
In the LOAM module, raw LiDAR point clouds are first continuously received by the system. These point clouds then undergo a pre-processing component to remove dynamic points from the vessel itself and to project the data onto a range image, which is a mapping in the angular coordinate, for segmentation. Next, an improved feature extraction method identifies distinct geometric features—such as surface, flat, and rough points. These features are then employed by a two-stage LiDAR Odometry component for pose estimation, which is processed further by a dedicated Plane Fitting component. Finally, the LiDAR mapping component incrementally builds the map and optimizes the trajectory by registering new scans against the global representation, integrating pose graph optimization via incremental smoothing and mapping (iSAM2) [41] within the Georgia Tech Smoothing and Mapping (GTSAM) [42] library to ensure global consistency.
Figure 2: Overview of the Inland-LOAM framework, which processes LiDAR data to generate IENC-compatible structural semantic maps for inland waterways. The LOAM pipeline (upper section) performs robust LiDAR odometry and constructs a globally consistent map. The semantic interpretation and conversion module (lower section) then uses this map to build a Voxel Map, analyze geometric properties for structural semantic understanding, and produce a 2D map converted into an IENC-compatible format.
The output from the LOAM module, comprising the point cloud map and estimated poses, serves as the input for the Conversion module in Fig. 2 lower part. In this module, a Voxel Map is created by dynamically combining features from the LOAM output based on probabilistic occupancy grids [43]. Next, 3D free and occupied space are identified and extracted from this Voxel Map. Slope estimation is then performed based on the extracted occupied regions. These estimated slopes, combined with local neighbourhood features from the Voxel Map, are subsequently used to determine structural semantic attributes, which collectively form a 2D Structural Semantic Map. Finally, a last conversion stage converts this 2D map into lightweight data that can be integrated with standard navigational charts,
3.2 System Equipment
The data acquisition setup, as illustrated in Fig. 3, integrates a Robosense Helios-32 LiDAR with a Septentrio AsteRx-i3 D Pro+ GNSS/INS receiver.
Figure 3: The layout of sensors on our self-developed SensorBox.
In response to the inherent water-surface interference challenges of inland waterways discussed in Section 2.1 [24, 25], Allis et al. [44] demonstrated that optimal detection of water surfaces relies on diffuse reflectivity, influenced predominantly by laser incidence angles and water turbidity. Furthermore, optical constraints, such as grazing angles, inherently limit the detectable surface area. Therefore, a LiDAR with an expanded vertical field-of-view was selected to maximize the reliability of water surface detection.
The Helios-32 is 32-line LiDAR with a maximum measurement range of 150 m and an accuracy of ± 1 cm. It provides a 360 degree horizontal field of view and a 70 degree vertical FoV, characterized by a non-uniform distribution of its scan lines and an approximate vertical angular resolution of 1.33 degrees. In this paper, the LiDAR was operated at a scanning frequency of 10 Hz and a horizontal angular resolution of 0.2 degrees. The AsteRx-i3 D Pro+ GNSS/INS receiver, equipped with dual antennas and an integrated Inertial Measurement Unit (IMU), provides acceleration and angular velocity measurements. It employs an internal Kalman filter to fuse IMU and GNSS data, delivering reliable centimeter level positioning and 3D orientation at an output frequency of 20 Hz for our experiments.
The intrinsic parameters of the IMU were calibrated using Allan Variance [45]. Extrinsic parameters calibration between the LiDAR and IMU was determined using GRIL-Calib [46]. Data processing and storage are managed by a laptop configured with an AMD Ryzen 9 PRO 7940HS CPU and 64 GB of RAM operating on Ubuntu 22.04, ROS 2111https://www.ros.org/.
3.3 LiDAR Odometry And Mapping Module
3.3.1 Preprocessing and Plane Fitting
The preprocessing begins by converting the raw LiDAR measurements into a structured range image. Let $\mathbf{P}_{k}=\{\mathbf{p}_{i}\}_{i=1}^{N}$ represent the LiDAR point cloud collected in the $k$ th scan cycle, where each point $\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]^{\top}\in\mathbb{R}^{3}$ denotes spatial coordinates. Due to the substantial vertical FoV of the LiDAR and the considerable physical dimensions of the vessel, some of the LiDAR returns inevitably originate from the vessel itself. Therefore, a bounding-box ($\text{bbox}_{\text{vessel}}$) filtering step is applied to exclude these points:
$\mathbf{P}_{k}^{\text{valid}}=\{\mathbf{p}_{i}\mid\mathbf{p}_{i}\notin\text{bbox}_{\text{vessel}}\}.$ (1)
Given the LiDAR’s non-uniform vertical distribution of its 32 laser beams and a fixed horizontal resolution of $0.2^{\circ}$, each scanning cycle produces a 2D range image, $I_{k}$, of dimension $32\times 1800$ (rows $\times$ columns). Each row corresponds to an individual laser beam, $v_{j}$ represents the scan line index (row), while each column represents sequential horizontal angle increments, $u_{j}$ represents the horizontal angle index (column). The resulting range image is shown in Fig. 4.
Figure 4: The single-frame point cloud is preprocessed with the vessel docked alongside the shore. In the upper figure, different laser beams are colored according to their vertical angles, and the red bounding box indicates the physical dimensions of the vessel. Several scan lines are observed to intersect the vessel, and these points are subsequently identified and removed. The lower figure shows the result after preprocessing. The water surface points are represented by blue, and the non-water surface points by yellow.
For subsequent segmentation tasks, including water surface point separation, outlier removal, and clustering of non-surface points, the data is arranged into an information matrix representation of dimensions $32\times 1800\times 3$. Specifically, the three channels are configured as follows:
  • 1.
    Channel 1 (Spatial Coordinates): ${P}_{k}^{\text{valid}}$ stores the 3D Euclidean coordinates $(x,y,z)$ of each valid pixel ($v_{j}$, $u_{j}$) in $I_{k}$.
  • 2.
    Channel 2 (Water Surface Indicator): $M_{k}^{surface}(v_{j},u_{j})$ is a binary classification result from water surface point segmentation, marking each point as surface or non-surface.
  • 3.
    Channel 3 (Cluster Labels): $M_{k}^{label}(v_{j},u_{j})$ contains the identifiers of clusters; isolated points classified as outliers are specifically labeled.
Our system is designed for short-range localization and mapping, the curvature of the Earth can be neglected, and the water surface can be approximated as a plane [47]. Moreover, considering the limitations discussed above posed by the incidence angles of grazing on the surface of the water, the candidate points for the estimation of the plane are restricted to a predefined subset of lower LiDAR beams ($N_{\text{lines}}$), determined by the installation height of the LiDAR installation and the draft of the vessel.
Unlike the commonly used ground plane extraction methods  [48], which iteratively ground points based on vertical angle differences between adjacent beams. For water surfaces, local geometric consistency based on angular criteria is susceptible to disruption by wave-induced undulations. RANSAC [49] is then used to fit these candidate points to the water surface plane. The estimated water surface can be mathematically described as follows:
$\mathbf{n}_{s}^{\top}\mathbf{p}+d_{s}=0$ (2)
where $\mathbf{n}_{s}=[a_{s},b_{s},c_{s}]^{\top}\in S^{2}$ is the normal vector of the plane and $d_{s}\in\mathbb{R}$ is its scalar offset from the sensor origin. Subsequently, any point $\mathbf{p}_{j}\in\mathbf{P}_{k}^{valid}$ is classified as a water surface point, setting the corresponding $M_{k}^{ground}(v_{j},u_{j})=1$, if its perpendicular distance to this estimated plane, calculated as $|\mathbf{n}_{s}^{\top}\mathbf{p}_{j}+d_{s}|$, is within a predefined tolerance $\tau_{s}$.
The subsequent step is to segment all remaining non-surface points. The segmentation is adapted from a range image-based clustering method [50] operating on the range image $I_{k}$. For each unlabeled point, a Breadth-First Search (BFS) is initiated to find connected components. Connectivity between an active point $\mathbf{p}_{\text{current}}=\mathbf{I}_{k}(v,u)$ and an adjacent non-surface point $\mathbf{p}_{\text{neighbor}}=\mathbf{I}_{k}(v^{\prime},u^{\prime})$ is evaluated based on a geometric criterion. This criterion assesses the 3D spatial relationship between $\mathbf{p}_{\text{current}}$ and $\mathbf{p}_{\text{neighbor}}$ by considering their respective ranges and the angular separation between them, as specified by their pixel indices $(v,u)$ and $(v^{\prime},u^{\prime})$ and the known angular resolutions. Points are clustered into the same segment if there is no significant depth discontinuity or an overly sharp angle between them according to a threshold $\theta_{\text{seg}}$.
All points belonging to a connected component found via BFS are assigned a unique positive integer identifier in Channel 3 ($M_{k}^{\text{label}}$). To ensure the robustness of subsequent feature extraction, a final filtering step is applied: segments comprising fewer than 30 points are considered insignificant or noise. Points belonging to these small segments are relabeled in Channel 3 as outliers. Fig 4 illustrates the result after preprocessing.
3.3.2 Feature Extraction
Algorithm 1 Feature Extraction
0: scan: $\mathcal{S}$
0: feature point sets: edge $\mathcal{E}$, flat $\mathcal{P}$
0:  threshold: $\tau_{1}$, $\tau_{2}$, $\tau_{3}$
[…]
2:for all scan line $L_{k}$ in $\mathcal{S}$ do
3:  for all segment line $G_{m}$ in scan line $L_{k}$ do
4:   for all point $p_{i}$ in scan line $G_{m}$ do
5:    Select neighbouring points:
6:    $\mathcal{M}=\{\,p_{i-n},\ldots,p_{i},\ldots,p_{i+n}\}$
7:    Calculate Curvature $c_{i}$
8:    $c_{i}=[\;\frac{1}{2n}\sum_{\tilde{p}\in\mathcal{M}}(p_{i}-\tilde{p})]^{2}$
9:    Calculate Roughness $r_{i}$
10:    $\bar{p}=\;\frac{1}{2n+1}\sum_{\tilde{p}\in\mathcal{M}}\tilde{p}$
11:    $r_{i}=\;\frac{1}{2n+1}\sum_{\tilde{p}\in\mathcal{M}}(\tilde{p}-\bar{p})^{2}$
12:    if $c_{i}>\tau_{1}\;\lor\;r_{i}>\tau_{2}$ then
13:     $\mathcal{E}\leftarrow\mathcal{E}\cup\{p_{i}\}$
14:    else if $r_{i}<\tau_{3}$ then
15:     $\mathcal{P}\leftarrow\mathcal{P}\cup\{p_{i}\}$
16:    end if
17:   end for
18:  end for
19:end for
20: Exclude unstable points from $\mathcal{E}$, $\mathcal{P}$ caused by occlusion
21: Non‐maximum Suppression on $\mathcal{E}$, $\mathcal{P}$
22:return  $\mathcal{E},\mathcal{P}$
Our feature extraction process is similar to the method in [15], but with optimized roughness characterisation for structural semantic mapping. Features are extracted from the segmented ground planes and surfaces identified in the preceding clustering step. A key challenge involves the robust discrimination of surface texture characteristics at local regions.
Traditional approaches often extract features based on local range differences. For instance, a common metric (Equation 3, where $|S|$ is the size of the local neighborhood, e.g., 10 points, and $r_{i}$ is the range of point $i$) is:
$c=\frac{1}{|S|\cdot\|r_{i}\|}\left\|\sum_{j\in S,j\neq i}(r_{j}-r_{i})\right\|$ (3)
While this method is computationally efficient and applicable to all directions, it demonstrates limitations in inland waterway environments. In areas with dense vegetation, grass, or gravel, the local surface exhibits numerous small, “edge-like” bumps. When applying Equation 3, the positive and negative range residuals $(r_{j}-r_{i})$ from neighboring points $p_{j}$ on either side of the central point $p_{i}$ tend to cancel each other out. Consequently, such curvature estimations often overlook subtle surface textures and primarily respond only to more extreme geometric features like significant peaks or occlusions.
To overcome this limitation and effectively capture local surface roughness, thereby distinguishing between smooth and textured surfaces, we employ a variance-based approach to quantify local range variations as detailed in Algorithm  1. Fig. 5 illustrates the comparative performance of feature extraction between LOAM-based methods and our optimized approach.
Figure 5: Comparative performance of feature extraction. Purple points indicate planar features, while green points denote rough features. The upper figure illustrates the feature distribution by our method, where semantic information of the targets is annotated, including building facades, trees, shore, and water surface. Notably, the shore is largely covered by vegetation, with only a small portion appearing as a flat surface. The lower figure shows the feature distribution of LOAM-based methods.
3.3.3 LiDAR Odometry, Mapping and Joint Optimization
Our LiDAR odometry and mapping parts build upon LEGO-LOAM [15], with specific adaptations in Joint Optimization. The system employs a two-stage registration process for pose estimation, integrating water surface constraints. The overall factor graph structure is illustrated in Fig. 6.
Two-Stage LiDAR Odometry: a two-step Levenberg-Marquardt optimization to estimate the transformation between consecutive scans. First, the planar features $\mathcal{F}_{p}^{t}$ extracted from the water surface and other planar environmental elements are matched to their correspondences in $\mathcal{F}_{p}^{t-1}$ to determine $t_{z},\theta_{roll},\theta_{pitch}]$. Then, the remaining transformation parameters $t_{x},t_{y},\theta_{yaw}]$ are estimated using edge features $\mathcal{F}_{e}^{t}$ matched against $\mathcal{F}_{e}^{t-1}$, while applying the previously computed vertical and angular parameters as constraints, resulting in the full 6 DoF transformation. This decoupled registration method significantly reduces computation time and is well-suited to our edge computing scenario.
LiDAR Mapping: To refine pose estimation, the LiDAR mapping module matches the extracted features $\{\mathcal{F}_{e}^{t},\mathcal{F}_{p}^{t}\}$ to a surrounding point cloud map at a lower frequency than the odometry module. Unlike the original LOAM that maintains a single point cloud map, we store individual feature sets $\{\mathcal{F}_{e}^{t},\mathcal{F}_{p}^{t}\}$ along with their corresponding poses. The surrounding map is constructed by transforming and fusing feature sets whose poses are within the sensor’s field of view (typically within 100m of the current position). This strategy of storing discrete feature sets with their poses facilitates flexible factor graph optimization. The resulting odometry estimates provide a continuous trajectory of the vessel. However, as with any LiDAR-based odometry, these estimates accumulate drift over time, particularly along the vertical axis due to the typical scarcity of distinct, stable horizontal features in many inland waterway environments.
Joint Optimization with Water Surface Constraints: To mitigate drift, we extend the pose graph optimization framework to incorporate global water surface constraints. We parameterize the water surface as a globally consistent plane $P_{w}$ with parameters $\{\mathbf{n}_{w},d_{w}\}$ as defined in Equation 2.
Figure 6: Inland-LOAM factor graph structure, showing prior, water surface, LiDAR odometry, loop closure factors.
The objective of our estimation is to determine the optimal set of robot poses $\mathcal{T}=\{T_{k}\}_{k=0}^{N}$, where $T_{k}\in SE(3)$. This is achieved by minimizing a cost function composed of several weighted squared error terms, representing the negative log-likelihood of the measurements given the states and plane parameters. The optimization problem is formulated as finding $\mathcal{X}^{*}=\{\mathcal{T}^{*},P_{w}^{*}\}$ by minimizing:
$\begin{split}\mathcal{X}^{*}=\arg\min_{\mathcal{T},P_{w}}\;&\|r_{0}\|^{2}_{\Sigma_{0}}\\ &+\sum_{k}\bigl{\|}r_{\mathrm{odom},k}(T_{k},T_{k+1})\bigr{\|}^{2}_{\Sigma_{\mathrm{odom},k}}\\ &+\sum_{(i,j)\in\mathcal{L}}\bigl{\|}r_{\mathrm{loop},ij}(T_{i},T_{j})\bigr{\|}^{2}_{\Sigma_{\mathrm{loop},ij}}\\ &+\sum_{m\in\mathcal{M}}\bigl{\|}r_{\mathrm{plane},m}(T_{m},P_{w})\bigr{\|}^{2}_{\Sigma_{\mathrm{plane},m}}\end{split}$ (4)
In this formulation, $\|r\|^{2}_{\Sigma}$ denotes the squared Mahalanobis distance $r^{T}\Sigma^{-1}r$, where $\Sigma$ is the covariance matrix of the corresponding residual $r$. The term $r_{0}$ represents a combined prior residual on initial knowledge about the system, such as the initial pose $T_{0}$ or an initial estimate for the plane, with $\Sigma_{0}$ as its covariance. The odometry residual, $r_{\mathrm{odom},k}$, for the relative motion $\Delta\widetilde{T}_{k,k+1}$ measured between consecutive poses $T_{k}$ and $T_{k+1}$, is given by
$r_{\mathrm{odom},k}=\mathrm{Log}\bigl{(}(\Delta\widetilde{T}_{k,k+1})^{-1}\,(T_{k}^{-1}T_{k+1})\bigr{)}\in\mathbb{R}^{6},$ (5)
where $\mathrm{Log}(\cdot)$ maps an $SE(3)$ element to its tangent space representation. Similarly, the loop closure residual, $r_{\mathrm{loop},ij}$, for a measured relative transform $\Delta\widetilde{T}_{i,j}$ between non-consecutive poses $T_{i}$ and $T_{j}$ (where $(i,j)\in\mathcal{L}$ denotes the set of loop closures), is defined as
$r_{\mathrm{loop},ij}=\mathrm{Log}\bigl{(}(\Delta\widetilde{T}_{i,j})^{-1}\,(T_{i}^{-1}T_{j})\bigr{)}\in\mathbb{R}^{6}.$ (6)
The sum $\sum_{m\in\mathcal{M}}$ accounts for all water plane observations, where $\mathcal{M}$ is the set of keyframes with such observations.
The water plane residual, $r_{\mathrm{plane},m}(T_{m},P_{w})$, arises from observing the water surface at a keyframe with pose $T_{m}=(R_{m},t_{m})$. A local plane $P_{s,m}=(n_{s,m},d_{s,m})$ is measured in the sensor frame, satisfying
$n_{s,m}^{T}x_{s}+d_{s,m}=0.$ (7)
This locally observed plane is transformed into the global coordinate frame by $T_{m}$ to obtain a predicted global plane observation $\widehat{P}_{w,m}=(\hat{n}_{w,m},\hat{d}_{w,m})$. The predicted global normal is
$\hat{n}_{w,m}=R_{m}n_{s,m},$ (8)
and the predicted global offset is derived from the transformed plane equation
$(R_{m}n_{s,m})^{T}x_{w}+\bigl{(}d_{s,m}-(R_{m}n_{s,m})^{T}t_{m}\bigr{)}=0,$ (9)
which gives
$\hat{d}_{w,m}=d_{s,m}-(R_{m}n_{s,m})^{T}t_{m}.$ (10)
[…]
$\begin{split}r_{\mathrm{plane},m}(T_{m},P_{w})&=P_{w}\ominus\widehat{P}_{w,m}\\ &=\begin{pmatrix}e_{N}\!\bigl{(}n_{w},\hat{n}_{w,m}\bigr{)}\\ d_{w}-\hat{d}_{w,m}\end{pmatrix}\\ &=\begin{pmatrix}e_{N}\!\bigl{(}n_{w},R_{m}n_{s,m}\bigr{)}\\ d_{w}-\bigl{(}d_{s,m}-(R_{m}n_{s,m})^{\mskip 2.0mu\top}t_{m}\bigr{)}\end{pmatrix}\in\mathbb{R}^{3}.\end{split}$ (11)
Here, $e_{N}(n_{w},\hat{n}_{w,m})\in\mathbb{R}^{2}$ is the 2D error vector representing the difference in orientation between the global plane normal $n_{w}$ and the predicted global normal $\hat{n}_{w,m}$ from the measurement. This error is typically computed by first defining a 3D error vector $\hat{\xi}$ in the tangent space of $n_{w}$, defined as follows:
$\hat{\xi}=-\frac{\arccos(n_{w}^{T}\hat{n}_{w,m})}{\sqrt{1-(n_{w}^{T}\hat{n}_{w,m})^{2}}}\,(\hat{n}_{w,m}-(n_{w}^{T}\hat{n}_{w,m})\,n_{w}),$ (12)
Then,the error $e_{N}$ is then obtained by projecting $\hat{\xi}$ using the basis $B_{n_{w}}$:
$e_{N}(n_{w},\hat{n}_{w,m})=B_{n_{w}}^{T}\hat{\xi}$. (13)
Here, $B_{n_{w}}\in\mathbb{R}^{3\times 2}$ is a matrix whose columns form an orthonormal basis for the 2D tangent space at the point $n_{w}$ on the unit sphere $S^{2}$.
By leveraging the water surface as a global geometric constraint, our approach reduces vertical drift without requiring external references such as GNSS. Fig. 7 demonstrates how this approach effectively maintains a consistent water surface representation throughout the trajectory.
Figure 7: The left image presents the point cloud map generated by our method in this scenario, while the right image shows the corresponding camera image captured in the same scene.
3.4 Structural Semantic Map Conversion
3.4.1 Semantic voxel transformation
In the preceding section, a map comprising structured point cloud features was generated. Each point in this cloud has been classified with semantic labels, including water surface, flat points, or rough points. To efficiently store and query this semantic information within the global map, the ”dense” semantically annotated point cloud is transformed into a ”sparse” voxel grid representation. For each voxel, a multi-class semantic confidence accumulation is maintained, based on the principles of probabilistic occupancy grids [43].
The workspace is discretized into a 3D voxel grid with a uniform resolution $r$, set to $0.15\text{ m}$ in this work. This resolution is chosen considering a balance between memory consumption, computational efficiency, and the typical scale of objects of interest in inland waterways. The entire transformation process is depicted in Fig. 8. For each voxel $v_{i}$, we maintain a hash map to store its semantic state $S(v_{i})$. This state involves accumulating evidence for the semantic classes: water surface, flat terrain, and rough terrain. The ’unknown’ category is assigned if a voxel has insufficient evidence for any specific class.
Figure 8: Overview of the Structural Semantic Map Conversion. Red voxels represent non-water voxels, while blue voxels indicate water voxels. During the transformation process, rough and planar points are accumulated as log-odds values and ultimately classified into three semantic categories: flat, rough, or unknown. Water voxels directly inherit the water surface points from the point cloud. The voxel map further supports slope estimation and bridge clearance calculation. For slope estimation, the yellow voxel represents the target cell, and a square neighborhood window is used for local slope computation. Bridge clearance is determined by measuring the vertical distance from the water surface voxels to the nearest non-water voxels above.
Points pre-classified as water surface by the LiDAR odometry module are handled directly, if such a point falls into voxel $v_{i}$, the voxel is immediately assigned the ’water’ semantic label with high confidence.
For terrain characteristics, we maintain log-odds values within the hash map entry for each non-water voxel. The initial belief for each of these log-odds components in every newly considered voxel is denoted as $l_{0,i}$.
When a new semantic point $p_{t}$ observed at time $t$, belonging to a terrain class $c\in\{\text{flat, rough}\}$, falls within voxel $v_{i}$, the log-odds for that class in the voxel is updated as follows:
$l_{t,i}^{c}=\text{logit}(p(m^{c}|y_{t}))+l_{t-1,i}^{c}-l_{0,i}^{c}$ (14)
where the log-odds in voxel $v_{i}$ at time $t$ is $l_{t,i}^{c}$. The term $p(m^{c}|y_{t})$ represents the probability of observing class $c$ given the measurement $y_{t}$. The logit function is defined as $\text{logit}(p)=\ln\left(\frac{p}{1-p}\right)$. We define the observation model probabilities for semantic classes, $p(\text{flat})=p_{f}$ (e.g., $p_{f}=0.7$) and $p(\text{rough})=1-p_{f}$.
Following the update, the log-odds $l_{i}^{c}$ for each voxel $v_{i}$ and each class $c$ are converted back to probabilities:
$P_{i}^{c}=\sigma(l_{i}^{c})=\frac{1}{1+e^{-l_{i}^{c}}}$ (15)
Ultimately, each voxel $v_{i}$ is assigned a definitive semantic label. Based on the probability, $P_{i}^{c}$:
$S(v_{i})=\begin{cases}\text{'flat'}&\text{if }P_{i}^{\text{flat}}-0.5>\tau_{\text{flat}}\\ \text{'rough'}&\text{if }P_{i}^{\text{flat}}-0.5<-\tau_{\text{rough}}\\ \text{'unknown'}&\text{otherwise}\end{cases}$
3.4.2 2D structural semantic map construction
The 3D semantic voxel map is subsequently projected onto a unified 2D structural semantic map. While it refers on the conversion method [40], this variant targets USV rather than UGV/UAV scenarios and adopts different map architectures for refined slope estimation and richer semantic inference. To effectively map the complex environment encountered by USVs in inland waterways, this framework broadly classifies non water surface areas into four structural semantic categories. These are: Horizontal Planar Surfaces (HP), which include ground, level terrain, and flat artificial or natural shore platforms; Vertical Planar Surfaces (VP), encompassing the facades of structures like bridges, buildings, and vessels, as well as vertical embankments; Complex Horizontal Terrain (CH), referring to irregular embankments, grasslands and low-lying scrub; and Complex Vertical Structures (CV), representing non-planar, upright natural elements like trees and dense bankside vegetation. Additionally, two further categories are defined for features on and over the water: Open Water Surface (OW) and Under Bridge Clearance Zone (UB), which identify the safe, traversable area beneath bridge structures.
The foundation of this conversion is an intermediate 2D grid representation, where each cell $(x,y)$ retains detailed information regarding the vertical distribution of voxels and their semantic labels derived from the 3D voxel map. Specifically, for each cell $(x,y)$ at the defined map resolution, we consider the voxel set: $V_{xy}=\bigl{\{}(z_{j},\,s_{j},\,p^{\mathrm{orig}}_{j})\bigr{\}}_{j=1}^{N_{xy}}$ where $z_{j}$ is the voxel height, $s_{j}$ is the semantic label (e.g., water, flat, rough and unknow), $p^{\mathrm{orig}}_{j}$ denotes the original 3D coordinates, and $N_{xy}$ is the total number of voxels in the vertical column at (x, y). For outdoor environments lacking explicit ceilings, the local maximum voxel height $Z_{max}$ serves as an initial estimate for the upper boundary of free space above newly detected occupied regions.
The next step is to perform slope estimation on these voxel sets, as illustrated in Fig. 8. In contrast to methods that derive slope using only boundary heights, our approach calculates the slope through all voxel points within a defined local neighbourhood. This method is more suitable for estimating those structures such as VP, CV, or overhangs.
For a target cell $(x_{c},y_{c})$, we define a square neighborhood window $W_{xy}$ of size $(2S_{a}+1)\times(2S_{a}+1)$, where $S_{a}$ specifies the extent of the neighbourhood.
Each voxel in $W_{xy}$ is represented by its relative coordinates and height $(x_{k}^{\mathrm{rel}},\,y_{k}^{\mathrm{rel}},\,z_{k})$, measured with respect to $(x_{c},y_{c})$. Let $N$ be the total number of such voxel data points aggregated from all cells within the entire window $W_{xy}$. The slope estimation is then performed using a least squares fit of a plane. The plane fitting is formulated as:
$\underbrace{\begin{bmatrix}x^{\mathrm{rel}}_{1}&y^{\mathrm{rel}}_{1}&1\\ x^{\mathrm{rel}}_{2}&y^{\mathrm{rel}}_{2}&1\\ \vdots&\vdots&\vdots\\ x^{\mathrm{rel}}_{N}&y^{\mathrm{rel}}_{N}&1\end{bmatrix}}_{\textstyle A}\underbrace{\begin{bmatrix}a\\ b\\ c\end{bmatrix}}_{\textstyle\theta}=\underbrace{\begin{bmatrix}z_{1}\\ z_{2}\\ \vdots\\ z_{N}\end{bmatrix}}_{\textstyle B}.$ (16)
The parameters $\theta_{\mathrm{ls}}=[a,b,c]^{\top}$ defining the best-fit plane are determined by solving the linear least squares problem shown in Eq. (16). To ensure numerical stability, the solution is obtained by computing the Moore–Penrose pseudoinverse $A^{+}$ of $A$ via singular value decomposition (SVD), rather than by explicitly forming normal equations.
$\theta_{\mathrm{ls}}\;=\;A^{+}\,B,$ (17)
Subsequently, the slope $m_{xy}$ at the central cell $(x_{c},y_{c})$ is computed as:
$m_{xy}=\sqrt{\,a_{\mathrm{ls}}^{2}+b_{\mathrm{ls}}^{2}}.$ (18)
The structural label for each 2D grid cell is determined by comparing three cues—$N$ (the number of reliable voxels in the window); $R_{\mathrm{plane}}$ (the fraction of those voxels that belong to locally planar surfaces $R_{\mathrm{plane}}\;=\;\frac{N_{\mathrm{plane}}}{N_{\mathrm{total}}}$); and the local slope $m_{xy}$. Let $\{N_{\mathrm{low}},N_{\mathrm{high}}\}$, $\{R_{\mathrm{low}},R_{\mathrm{high}}\}$, and $\{m_{\mathrm{low}},m_{\mathrm{high}}\}$ denote a set of heuristic rules. A cell is classified as:
  • 1.
    HP (Horizontal Plane) if $N>N_{\mathrm{high}}$, $R_{\mathrm{plane}}>R_{\mathrm{high}}$ and $m_{xy}<m_{\mathrm{low}}$;
  • 2.
    VP (Vertical Plane) if $N>N_{\mathrm{low}}$, $R_{\mathrm{plane}}>R_{\mathrm{low}}$ and $m_{xy}>m_{\mathrm{high}}$;
  • 3.
    CH/CV (Complex Horizontal / Vertical) whenever the slope is appreciable ($m_{xy}\geq m_{\mathrm{low}}$) but the planar support is insufficient for VP or HP;
A grid cell that contains only water-labelled voxels is marked as Open Water (OW). If a second, non-water voxel layer appears more than a predefined safety margin $\tau_{\mathrm{bridge}}$ (e.g., $3.0\,\mathrm{m}$) above the water layer, the cell is flagged as Under-Bridge (UB) the vertical clearance, $G_{\mathrm{bridge}}$ is determined by the difference $G_{\mathrm{bridge}}=Z_{\text{overhead\_lowest}}-Z_{\text{water\_level}}$, where $Z_{\text{overhead\_lowest}}$ is the Z-axis of the lowest point of the overhead structure and $Z_{\text{water\_level}}$ represents the Z-axis of the highest point water surface within the vertical column. These additional rules override the geometric label whenever they trigger, ensuring bridges are never misclassified.
To provide a more practical representation of the clearance under the bridge, we first divide the 2D grid map into local blocks of size $l\times l$ and compute the mean and range of $G_{\mathrm{bridge}}$ within each block. These blocks are then further grouped into larger contiguous regions using BFS. For each region, the mean clearance of its constituent blocks is taken as the representative value, so that each under-bridge area is assigned a single vertical clearance. This two-stage aggregation ensures both the accuracy and real-time performance of the measurement.
3.4.3 Generation of IENC Compatible Map
This process transforms costmap data into a durable and shareable representation, aiming for compatibility with IENC standards. It involves extracting key features, such as shorelines and bridges, and representing them with distinct geometric primitives. The method begins by encoding specific occupancy categories into an RGB image where each category is assigned a unique color value. Subsequently, two primary types of features are extracted, detailed in Algorithm 2.
For Shoreline Extraction Represented by Points, a ray casting method is used to delineate shorelines. From the boundaries of identified water regions, rays are projected outwards in multiple directions. The end points of these rays, where they intersect the nearest non-water pixels, define the shoreline.
For Bridge Extraction Represented by Polygons, bridge pixels are first identified and then clustered using DBSCAN [51] to group fragmented detections into coherent structures. For each cluster, a convex hull is computed to generate a polygonal representation. Overlapping or closely spaced polygons are merged to ensure continuous bridge structures are represented as unified entities. The resulting polygons are then exported in JSON format for integration with navigation systems.
Algorithm 2 Costmap $\to$ IENC Feature Extraction
0: RGB image $\mathcal{I}$
0: Shoreline points $\mathcal{S}$, bridge polygons $\mathcal{B}$
0:  DBSCAN $(\varepsilon,\,n_{\min})$, hull merge distance $\delta$, ray step $s$, angle step $\Delta\theta$ // ———- Shoreline extraction ———-
1: Define $\mathcal{B}_{w}$: set of water boundary pixels in $\mathcal{I}$
2: Define $\mathcal{O}$: set of obstacle mask pixels in $\mathcal{I}$
3: Define $\mathcal{Q}$: set of shoreline hit
[…]
5:for all $\mathbf{p}_{0}\in\mathcal{B}_{w}$ do
6:  for $\theta=0$ to $360-\Delta\theta$ step $\Delta\theta$ do
7:   for $k=1$ to max_steps do
8:    $\mathbf{p}\leftarrow\mathbf{p}_{0}+k\cdot s\cdot(\cos\theta,\sin\theta)^{\top}$
9:    if  $\mathbf{p}$ is out of bounds  then
10:     break
11:    end if
12:    if  $\mathbf{p}$ is not water  then
13:     if  $\mathbf{p}$ is obstacle  then
14:      $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{\mathbf{p}\}$; break
15:     else
16:      break
17:     end if
18:    end if
19:   end for
20:  end for
21:end for
[…]
23: Define $\mathcal{P}_{\text{bridge}}$: set of bridge pixels in $\mathcal{I}$
[…]
25: Initialize set of bridge polygons
[…]
27:for  each cluster $C$ in labels  do
28:  $\mathcal{H}\leftarrow\mathcal{H}\cup\{\textsc{ConvexHull}(C)\}$
29:end for
30: Merge polygons within $\delta$ distance
[…]
32:return  $\mathcal{S},\mathcal{B}$
4 Experiments and Discussion
4.1 Data Collection
To validate the performance of the proposed Inland-LOAM, we conducted three representative experiments on November along the canal from Leuven to the Dijle (5053.8890N, 00442.4010E). The test routes span various conditions, including urban canyons, vegetated banks, mooring zones, and bridge crossings, providing a challenging and realistic testbed for evaluating the robustness and precision of our semantic mapping and localization system.
Data were collected using a catamaran named the Maverick, with an overall length of 6.10 m and a beam of 2.02 m, as illustrated in Fig. 1. Its draft varies from approximately 0.30 m (lightship) to 0.60 m (fully ballasted). The twin‐hull configuration provides inherent stability, beneficial for sensor data acquisition. The vessel was equipped with the sensor suite detailed in Section 3.2.
We recorded three datasets along this route, as shown in Fig. 9, to cover different typical inland waterway environments and to demonstrate the semantic mapping capability. These trajectories are named as follows, listed in Table 1:
  • 1.
    Mixed Area: Between two fixed bridges, with quay walls, stone revetments, vegetated banks, and exposed soil.
  • 2.
    Vegetation Area: Both banks are densely vegetated, featuring several mooring facilities and multiple docked vessels.
  • 3.
    Urban Port Area: There are large buildings on both sides of the Port, including a fixed bridge and a lifting bridge, with multiple mooring facilities and small boats docked.
Figure 9: Our data collection was conducted along the canal from Leuven to the Dijle. The collected trajectories are visualized as follows: Area A in red, Area B in yellow, and Area C in green.
Table 1: Our Inland Waterway Datasets
Experiment Distance [m] Duration [s] Loop Back Weather
A (Mixed Area) 424.5 561.7 Sunny
B (Vegetation Area) 583.5 356.8 Overcast
C (Urban Port Area) 474.9 399.9 Overcast
4.2 Evaluation of Trajectory and Point Cloud Mapping
To highlight the performance of our method, we conducted comparative experiments against several state-of-the-art methods: KISS-ICP, HDL-SLAM, LeGO-LOAM, and LOAM. To ensure a fair and consistent comparison, all methods were adapted to match the laser beam distribution specific to our dataset.
To assess the quality of the generated point cloud maps, we compared Inland-LOAM visually with LOAM, LeGO-LOAM, HDL-SLAM, and KISS-ICP within Area B, as illustrated in Fig. 10. The gradient coloring along the trajectory indicates elevation changes. Notably lacking bridges or large overhead structures, Area B presents a challenge for evaluating vertical localization accuracy.
At a global scale, LOAM and KISS-ICP show a water surface that curves downwards along the z-axis, while HDL-SLAM exhibits an upward curvature. In contrast, LeGO-LOAM displays only minor z-axis drift. Inland-LOAM maintained a relatively flat water surface, demonstrating superior mitigation of vertical drift. At a local scale, Fig. 10 details the mapping quality of the same building produced by each method, reflecting the precision of their horizontal registration. KISS-ICP, LOAM, and LeGO-LOAM fail to reconstruct the building’s windows, capturing only the main facade. While HDL-SLAM correctly identifies the windows, it fails to represent their proper alignment. Inland-LOAM, however, accurately reconstructs both the windows and their correct spatial arrangement, a clear benefit of our optimized feature extraction algorithm.
The accuracy of the trajectory estimation was quantitatively assessed by comparing Inland-LOAM with LOAM, LeGO-LOAM, KISS-ICP, and HDL-SLAM. Fig. 11 presents these comparative trajectories. We calculated the Root Mean Square Error (RMSE) and Standard Deviation (STD) between RTK-GNSS ground truth positions and the trajectories generated by these methods, as summarized in Table 2. It is important to note that, although RTK-GNSS data were treated as ground truth, GNSS signals can experience drift in certain inland waterway scenarios, particularly under bridges or in narrow urban sections in Figs. 11(a) (sections 1 and 3) and (c) (right side).
As shown in Fig. 11(b), the Inland-LOAM trajectory closely overlaps the ground truth of RTK-GNSS, while the other methods exhibit a noticeable gap. In Figs. 11(a) and  11(c), HDL-SLAM displayed a larger gap, while Inland-LOAM consistently maintained a minimal deviation from the ground truth. As shown in Table 2, Inland-LOAM achieves the lowest RMSE and standard deviation in each test case. Overall, the trajectory results indicate that although all methods performed similarly in general localization within inland waterways, Inland-LOAM achieved superior accuracy and robustness under challenging conditions. This high-precision trajectory estimation provides a solid foundation for subsequent semantic analysis and mapping tasks.
Figure 10: Mapping results of Inland-LOAM, KISS-ICP, HDL-SLAM, LeGO-LOAM, and LOAM in the vegetation area, with a building detail by each method in this scene. The gradient coloring in the Z-axis direction indicates elevation changes.
(a) Area A
(b) Area B
(c) Area C
Figure 11: Comparison of the trajectories of each algorithm in experiments. Sub-figures (a-2), (b), and the left of (c) provide magnified views for examination of local accuracy. Sub-figures (a-1), (a-3), and the right of (c) project the trajectories from Inland-LOAM (blue) and the ground truth (black) onto IENC to demonstrate accuracy in confined spaces.
Table 2: Algorithm trajectory accuracy evaluation.
Inland-LOAM HDL-SLAM KISS-ICP LeGO-LOAM LOAM
Seq RMSE STD RMSE STD RMSE STD RMSE STD RMSE STD
A 0.45 0.37 3.14 1.34 1.28 0.64 0.48 0.37 1.27 0.58
B 0.60 0.30 4.87 2.11 9.61 4.78 2.26 1.14 7.44 3.59
C 0.99 0.72 2.25 1.06 2.22 1.08 2.01 0.83 1.97 0.97
Avr 0.68 0.46 3.42 1.50 4.37 2.17 1.58 0.78 3.56 1.71
4.3 Structural Semantic Mapping Evaluation
(a) Area A
(a) Area B
(a) Area C
Figure 14: Structural semantic maps generated by Inland-LOAM in areas A, B, and C. Each scenario is presented with corresponding real-world photographs from multiple perspectives, annotated to explain key features. The color encoding corresponds to the semantic classes defined in Section 3.4.2: red for Horizontal Planar Surfaces (HP), pink for Vertical Planar Surfaces (VP), green for Complex Horizontal Terrain (CH), yellow for Complex Vertical Structures (CV), blue for Open Water Surface (OW), and black for Under Bridge Clearance Zone (UB). Note that this color scheme is used for visualization purposes only and does not represent the traditional cost definitions in a costmap. For (a) and (c), detailed semantic maps of bridge areas are included, with the identified clearance height explicitly labeled (H).
This section evaluates the quality of these semantic maps and discusses their practical utility. We assess the performance of our method by comparing the generated structural semantic maps with their real-world counterparts.
Fig. 14 presents the 2D structural semantic maps generated for our three experimental areas A, B, and C. Fig. 12(a) shows the performance along an inland waterway with highly diverse shoreline characteristics. The system accurately identifies navigable areas under bridges, providing real-time bridge clearance estimations (H-1, H-2, H-3), which are essential for the safe passage of large vessels and address the limitations of traditional static IENCs. Furthermore, it demonstrates the ability to differentiate shoreline constructions, distinguishing boat ramps classified as HP at c-1, shore platforms (HP) at e-3, containers classified as VP at d-2, and bridge piers (VP) at e-1. It also distinguishes vegetated banks classified as CH and CV at a-2, and piling identified as CV at c-2.
Fig. 13(a) illustrates a scenario dominated by dense reed vegetation. The map correctly captures the overall characteristics of the canal, categorizing large buildings at location b-1 as VP and most of the shoreline as complex terrain CH or CV, aligning well with real-world scenes. However, this scenario highlights limitations of purely geometric classification. Scattered small HP and VP regions appear within extensively classified vegetation areas (e.g., near c-2 and d-2). Such ambiguity likely results from LiDAR penetrating sparse foliage and capturing flat ground or mistaking dense, flat-topped vegetation as planar surfaces. Although this reveals the inherent challenges of using geometric methods in complex natural environments, it also demonstrates the ability to detect subtle variations in surface structure.
In Fig. 14(a), the system successfully identifies bridges and labels the navigable water beneath UB. It clearly classifies large building facades in b-3 as VP, and precisely outlines vessels (b-2, c-2) and revetments (b-1, c-1). However, movement of the lifting bridge (H-1) during data acquisition led to color mixing between the UB and VP areas, revealing limitations of our system in dynamic environments.
These semantic maps represent not only geometric reconstructions but also serve as costmaps directly usable for navigation planning. For USV decision-making, the OW regions represent low-cost navigable areas, while the UB areas indicate regions requiring height awareness. Combined HP and VP regions, such as docks, can be considered potential docking platforms. Continuous VP boundaries adjacent to OW can indicate vessel or revetment boundaries, whereas VP boundaries distant from OW can represent building facades for higher-level semantic localization tasks. The CH and CV regions represent high-cost obstacles or areas that require careful avoidance to prevent grounding or vegetation entanglement of engines.
4.4 Shoreline Boundaries Comparison with Reference Maps
To evaluate the effectiveness of the boundaries derived from structural semantic mapping, we performed a comparative analysis with three widely used reference maps: Google Earth, OpenStreetMap (OSM) [^2] https://www.openstreetmap.org, and IENCs. As illustrated in Figure 15, each reference map was created using data from a different period: Google Earth used aerial imagery from on or before June 16, 2023; OSM is based on a continuously updated database; and the IENC data was last modified on July 1, 2021.
In Area A (Fig. 15(a)), shoreline boundaries (red hits) align with the contours of artificial structures such as revetments and shoreline constructions, exhibiting strong correspondence with the IENC and Google Earth. In contrast, for vegetated banks, some discrepancies appear relative to the satellite imagery, reflecting seasonal variation in reeds or minor erosion not represented in the static maps. In Area B (Fig. 15(b)), our system accurately delineates boundaries, including vessels moored long-term along the canal. Area C (Fig. 15(c)) demonstrates a distinct case. Although the canal boundary aligns with all reference datasets, the layout of moored vessels within the port basin shows a notable divergence. The positions identified by Inland-LOAM are consistent with the latest updates in OSM and differ from the outdated layouts present in other map, reflecting Inland-LOAM’s potential to dynamically enhance and maintain up-to-date navigational charts.
(a) Area A
(b) Area B
(c) Area C
Figure 15: The shoreline boundaries generated by Inland-LOAM were projected onto the IENC, Google Earth, and OpenStreetMap datasets in areas A, B, and C, respectively.
4.5 Time analysis
Table 3 evaluates the real-time performance of the LOAM module. The front-end components, which include pre-processing, plane fitting, feature extraction, and LiDAR odometry, have a combined processing time of 18.3 ms. The back-end mapping module is 37.2 ms. Both timings are shorter than the data acquisition interval required by the LiDAR’s sampling frequency (10 Hz) and the map generation frequency (2 Hz), respectively, confirming the real-time performance of our LOAM module.
For the conversion module, the computational time required to transform the voxel map into the 2D structural semantic map depicted in Fig. 16. The voxel map update time scales with map size, and this time cost is particularly pronounced in open environments compared to enclosed scenes [40]. In our experiments, Area A is the only trajectory that includes loop closures, resulting in significantly lower computation times compared to areas B and C, which are focused on open area exploration. This difference arises because in loop-back scenarios, only the updated map segments require modification, while open-area exploration triggers broader map updates. Despite this, the median map conversion time for areas B and C is approximately 1s, with a maximum of 2.8s. These update times are sufficient within the requirements for real-time decision making and control on a USV driving at up to 3m/s, given a 150m LiDAR sensing range.
The IENC integration module, primarily designed for long-term maintenance and sharing of known environment maps, is executed offline. This approach ensures that the generated IENC-compatible data maintains high accuracy and stability without imposing additional computational cost during navigation tasks.
Table 3: Processing Time for the Components of LOAM Module.
Component Time (ms)
Pre-processing & Plane Fitting 7.8
Feature Extraction & LiDAR Odometry 10.5
Mapping 37.2
Figure 16: The computation time to update the 2D map generated from the voxel map under different exploration areas.
5 Conclusion
In this paper, we proposed Inland-LOAM, an integrated LiDAR-only SLAM framework specifically designed for robust odometry, mapping, and structural semantic map generation in challenging inland waterway environments. Addressing the limitations of conventional SLAM methods, our contributions include an enhanced feature extraction approach and a joint optimization method that uses the water surface as a globally consistent planar reference to effectively reduce drift. Furthermore, we detailed an end-to-end pipeline transforming dense 3D point clouds into structured, semantic 2D maps. The proposed voxel-based semantic segmentation and subsequent 2D mapping method allow real-time estimation of critical navigational parameters, such as bridge clearances, and provide precisely delineated shorelines compatible with IENC standards.
Through comprehensive experiments conducted on our collected dataset, Inland-LOAM demonstrated superior performance in terms of localization precision and semantic map quality compared to the state-of-the-art methods. Semantic maps provide detailed, up-to-date, and actionable insights, enhancing navigational safety and operational efficiency for autonomous vessels.
However, our method has certain limitations. Purely geometric-based classification exhibits ambiguity in regions with complex natural textures, such as densely vegetated areas, leading to occasional misclassification at a local scale. Additionally, the current static mapping approach faces challenges when encountering dynamic elements, as evidenced by the aliasing artifacts caused by moving lifting bridges. Reliable detection of the water surface itself can also be compromised in highly dynamic circumstances, such as in the presence of nearby moving vessels, which can degrade the effectiveness of our global planar constraint and introduce additional drift. This highlights potential vulnerabilities when operating in environments with significant dynamic infrastructure or dense vessel traffic.
Future research will focus on addressing these limitations. One promising direction is to enhance semantic classification accuracy by integrating additional sensor modalities, particularly RGB cameras. Leveraging color and texture information could help resolve semantic ambiguities among geometrically similar but semantically distinct surfaces. To improve robustness in dynamic environments, the framework could be extended with temporal analysis and multi-object tracking capabilities. Moreover, the generated semantic maps pave the way for higher-level autonomous applications, such as semantic-guided localization using stable landmarks like building facades or revetments, and path planning based on semantic costmaps.
6 CRediT authorship contribution statement
Zhongbi Luo: Conceptualization, Methodology, Investigation, Resources, Validation, Visualization, Formal analysis, Software, Data curation, Writing – original draft, Writing – review & editing. Yunjia Wang: Writing – review & editing, Data curation. Jan Swevers: Writing – review & editing. Peter Slaets: Project administration, Supervision, Funding acquisition, Writing – review & editing. Herman Bruyninckx: Conceptualization, Methodology, Visualization, Supervision, Writing – review & editing.
7 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
8 Acknowledgment
We would like to thank the subject editor and anonymous reviewers who provided insightful and constructive comments on the manuscript. This project has received funding from the European Union’s EU Framework Programme for Research and Innovation Horizon 2020 under Grant Agreement No. 955768 (ETN AUTOBarge). Project website: https://etn-autobarge.eu/.